What is the price of a $1,000 par 9 year, annual bond with a 8.1% coupon rate and a yield to maturity of 6.3%? (Show your answer to nearest cent with no comma. For example $1,378.565 is entered as 1378.57)
Answer #1

Annual coupon=1000*8.1%=81

Hence price of bond=Annual coupon*Present value of annuity factor(6.3%,9)+$1000*Present value of discounting factor(6.3%,9)

=81*6.71376257+1000*0.577032958

=1120.85(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=81[1-(1.063)^-9]/0.063

=81*6.71376257

2.Present value of discounting factor=1000/1.063^9

=1000*0.577032958